OpenStudy (anonymous):

What is the sum of a 6-term geometric series if the first term is 22 and the last term is 369,754?

OpenStudy (anonymous):

AND what is the sum of an 8 term geometric series id the first term is 15 and the last term is -4,199,040?

OpenStudy (anonymous):

What is the sum of a 6–term geometric series if the first term is 22 and the last term is 369,754? first term, a = 22 last term = 369,754 the general form of a geometric sequence is a , a r , a r ² , a r ³ , .......... , a r ⁿ .................>>>...where a is the first term, .................>>>...r is the common ratio, .................>>>...nth term = a r ^ ( n - 1 ) last term or 6th term = ar^5 = 369,754 ( 22 ) r^5 = 369,754 r^5 = 16,807 r = 7 ≈≈≈≈≈≈ sum of the 6-term geometric series is: n Σ a r^k = a ( 1 - r^(n+1)) / ( 1 - r ) k=0 5 Σ 22 ( 7 )^k = 22 ( 1 - ( 7 )^(5+1)) / ( 1 - 7 ) k=0 ....................= 22 ( 1 - ( 7 )^6 ) / –6 ....................= 22 ( 1 - 117,649 ) / –6 ....................= 22 ( –117,648 ) / –6 ....................= –2,588,256 / –6 ....................= 431,376

OpenStudy (anonymous):

Did that help?

OpenStudy (anonymous):

oh my goodness, this is amazing! Thank you so much.

OpenStudy (anonymous):

You should message me. :)

OpenStudy (anonymous):

Please.

OpenStudy (anonymous):

will do!